# Classification of generalized Yamabe solitons

### v4

@@ 3. Gradient conformal solitons with Vanishing condition on Cotton and Cao-Chen
Roughly speaking, on gradient Ricci solitons, the Cao-Chen tensor estimates the difference between the Weyl tensor and the Cotton tensor, but, on gradient Yamabe solitons, it doesn’t. Therefore it is interesting to consider the gradient conformal solitons with $D\equiv 0$.
@@
A nontrivial complete gradient conformal soliton $(M^{n},g,F,\varphi)$ with vanishing Cao-Chen tensor is either
@@
where, $\bar{g}_{S}$ is the round metric on $\mathbb{S}^{n-1},$ or
@@
where, $N_{Ein}$ is an Einstein manifold. Furthermore, if $R\geq 0$, then either $\bar{R}>0$, or $R=\bar{R}=0$ and $(M,g)$ is isometric to the Riemannian product $(\mathbb{R},dr^{2})\times(N^{n-1},\bar{g})$, where, $N$ is Ricci flat.
@@
We only have to consider (3) of Theorem 1.1. By the same argument as in the proof of Theorem 1.1, one has
@@
Since the Cao-Chen tensor vanishes, and $F^{\prime}>0$, one has
@@
Hence, $N$ is an Einstein manifold. Thus, the scalar curvature $\bar{R}$ of $N$ is constant. Assume that $R\geq 0$. If $\bar{R}\leq 0$, then by (3.7), $F^{\prime\prime\prime}\leq 0$. Therefore, $F^{\prime}(>0)$ is concave, which means that $F^{\prime}$ is constant. By (3.7) again, $R=\bar{R}=0$ and $\overline{\rm Ric}=0$ on $N$. Therefore, $(M,g)$ is the Riemannian product $(\mathbb{R},dr^{2})\times(N^{n-1},\bar{g}),$ where, $N$ is Ricci flat. ∎
As is well known, Einstein manifolds of dimension $n\leq 3$ are space forms, hence one can get the following.
Corollary 3.4.
A nontrivial complete gradient conformal soliton $(M^{n},g,F,\varphi)$ $(n\leq 4)$ with vanishing Cao-Chen tensor is either
@@
where, $\bar{g}_{S}$ is the round metric on $\mathbb{S}^{n-1},$ or
@@
where, $(N^{n-1}(c),\bar{g})$ is a space form.
Therefore, we have the following.
Corollary 3.5.
Any complete nontrivial nonflat gradient conformal soliton $(M^{n},g,F,\varphi)$ $(n\leq 4)$ with $D\equiv 0$ and $R\geq 0$ is rotationally symmetric.
4. Classification of locally conformally flat gradient conformal solitons
In this section, inspired by [4], [6], [7] and [9], we classify locally conformally flat gradient conformal solitons.
Lemma 4.1.
Let $(M,g,F,\varphi)$ be a nontrivial complete locally conformally flat gradient conformal soliton. Assume that $F$ has no critical point. Then, $(M,g,F,\varphi)$ is warped product
$(\mathbb{R},dr^{2})\times_{|\nabla F|}\left(N^{n-1}(c),\bar{g}\right),$
where, $(N^{n-1}(c),\bar{g})$ is a space form.
Proof.
We consider $(3)$ of Theorem 1.1. By the same argument as in the proof of Theorem 1.1, one can get formulas of the warped product manifold of the warping function $(0<)|\nabla F|=F^{\prime}(r)$. For $a,b,c,d=2,3,\cdots n.$
Case 1. dim $M=3$: By Proposition 3.2, $N$ is a space form.
Case 2. dim $M\geq 4$: By (3.1), (4.1), (4.2) and (4.3), one has
$W_{1a1b}=$ $-\frac{\bar{R}_{ab}}{n-2}+\frac{\bar{R}}{(n-1)(n-2)}\bar{g}_{ab},$
$W_{1abc}=$ $0,$
$W_{abcd}=$ $(F^{\prime})^{2}\Big{(}\bar{W}_{abcd}$
Since $M$ is locally conformally flat, one has
(4.4) $\bar{R}_{ab}=\frac{\bar{R}}{n-1}\bar{g}_{ab},$
and
Substituting (4.4) into (4.5), one has $\bar{W}_{abcd}=0.$ Therefore, $N$ is Einstein and locally conformally flat, which means that $N$ is a space form. ∎
Combining Lemma 4.1 with Theorem 1.1, we obtain the following.
Corollary 4.2.
A nontrivial complete locally conformally flat gradient conformal soliton $(M^{n},g,F,\varphi)$ is either
$(1)$ compact and rotationally symmetric, or
$(2)$ rotationally symmetric and equal to the warped product
$([0,\infty),dr^{2})\times_{|\nabla F|}(\mathbb{S}^{n-1},{\bar{g}}_{S}),$
where, $\bar{g}_{S}$ is the round metric on $\mathbb{S}^{n-1},$ or
$(3)$ the warped product
$(\mathbb{R},dr^{2})\times_{|\nabla F|}\left(N^{n-1}(c),\bar{g}\right),$
where, $(N^{n-1}(c),\bar{g})$ is a space form.
Therefore, we have the following.
Corollary 4.3.
Any complete nontrivial nonflat locally conformally flat gradient conformal soliton with $R\geq 0$ is rotationally symmetric.
Data availability statement
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Conflict of interest
There is no conflict of interest in the manuscript.

### v5

@@ 4. Gradient conformal solitons with a vanishing condition on the Cao-Chen tensor
Roughly speaking, on gradient Ricci solitons, the Cao-Chen tensor estimates the difference between the Weyl tensor and the Cotton tensor, but an analogous interpretation does not hold for conformal gradient solitons.
Therefore, it is interesting to consider conformal gradient solitons with $D\equiv 0$.
@@
A nontrivial complete conformal gradient soliton $(M^{n},g,F,\varphi)$ with $D\equiv 0$ is either
@@
where $\bar{g}_{S}$ is the round metric on $\mathbb{S}^{n-1},$ or
@@
where $N_{Ein}$ is an Einstein manifold. Furthermore, if $R\geq 0$, then either $\bar{R}>0$, or $R=\bar{R}=0$ and $(M,g)$ is isometric to the Riemannian product $(\mathbb{R},dr^{2})\times(N^{n-1},\bar{g})$, where $N$ is Ricci flat.
@@
We only have to consider (3) of Lemma 2.1. By the same argument as in the proof of Lemma 2.1, we have
@@
Since the Cao-Chen tensor vanishes and $F^{\prime}>0$, we have
@@
Hence, $N$ is an Einstein manifold. Thus, the scalar curvature $\bar{R}$ of $N$ is constant. Assume that $R\geq 0$. If $\bar{R}\leq 0$, then by (4.2), $F^{\prime\prime\prime}\leq 0$. Therefore, $F^{\prime}(>0)$ is concave, which means that $F^{\prime}$ is constant. By (4.2) again, $R=\bar{R}=0$ and $\overline{\rm Ric}=0$ on $N$. Therefore, $(M,g)$ is the Riemannian product $(\mathbb{R},dr^{2})\times(N^{n-1},\bar{g}),$ where $N$ is Ricci flat. ∎
If $\dim N=n-1\leq 3$, then the Einstein manifold $N$ is a space form. Hence, we obtain the following theorem.
Theorem 4.2.
For $n=3,4$, a nontrivial complete conformal gradient soliton $(M^{n},g,F,\varphi)$ with $D\equiv 0$ is either
@@
where $\bar{g}_{S}$ is the round metric on $\mathbb{S}^{n-1},$ or
@@
where $(N^{n-1}(c),\bar{g})$ is a space form.
By the same argument as in Section 3, we obtain Theorem 1.2.